# Any-Order Flexible Length Masked Diffusion

### v1

@@ 3 Variable Length Masked Diffusions: Training
$s_{t}\colon=\{i\in\{1,\dots,\mathrm{len}(x_{1})\}\mid T_{1}^{i}\leq t\},$
@@
$f_{\theta}$ resembles the familiar unmasking posterior from MDMs, whereas $g_{\theta}$ is new: it predicts how many tokens need to be inserted. Intuitively, modeling a variable-length $p_{1}$ is harder than the fixed-length setup of MDM–introducing an insertion expectation allows us to parameterize more complicated CTMC for FlexMDM; its rate matrix will appear soon in Proposition 2. To define the training loss, we set the boundary values of $s_{t}$ as $s_{t}[-1]:=0$ and $s_{t}[\mathrm{len}(s_{t})]:=\mathrm{len}(x_{1})$, and let $\phi(x,y)=y-x\log y$ denote a scalar Bregman divergence.
@@ 5.1.1 Pretraining on text data
We first construct a training dataset from the raw OpenWebText corpus (Gokaslan et al., 2019), spliting each article into paragraphs to preserve semantic coherence and yield variable-length sequences. Models pretrained on this data therefore generate variable-length text.
Results. We train 175175M FlexMDM and MDM with a maximum sequence length $1024$ for 500500K iterations and batch size $1024$. Using the pretrained models, we vary the number of sampling steps and measure (a) generative perplexity as a proxy for text fluency, and (b) the induced length distribution. Figure 5(c) shows comparative generative perplexity for the two models, improving as sampling steps increase, indicating no fluency degradation for FlexMDM despite its more involved loss objective. Crucially, we observe that FlexMDM matches the true length distribution far more closely (Figure 5(b)): with only $256$ steps it tracks the ground truth distribution, whereas MDM remains miscalibrated even at $1024$ steps.
@@ 5.1.2 Planning task
We further evaluate FlexMDM’s ability in a planning task in a discrete space. Motivated by an earlier study Janner et al. (2022) that investigated the ability of continuous diffusion in maze tasks, we design a grid-maze benchmark: the maze is fixed but unknown to the model, with a subset of cells invalid. Given a sequence of subgoal grids $(g_{1},\dots,g_{K})$, the model must connect this sequence without entering invalid cells (see Figure 5(a)). This subgoal structure aligns naturally with FlexMDM: starting from $(g_{1},\dots,g_{K})$, inference inserts mask tokens between subgoals and then unmasks to generate a feasible path. Theoretically, this can be seen as augmenting the base distribution to contain data-dependent distribution Albergo et al. (2024). This is in stark constract to MDM, where it must preassign each subgoal to a specific position, which is difficult to know a priori. We provide additional details in Appendix F.2.
@@ 5.2 Scaling up FlexMDM
In this section, we address FlexMDM’s scalability by scaling it to 8B parameters and observing notable improvements over an MDM baseline. We start from the observation that MDM and FlexMDM both share the unmasking posterior as a core component, suggesting effective task transfer from a pretrained MDM might be possible. To demonstrate this, concretely, we initialize from LLaDA-Base (Nie et al., 2025) and make the following modifications: (a) add time-embedding layers and a scalar head to model the insertion expectation; (b) attach LoRA adapters. Altogether, the resulting number of trainable parameters is $\approx$400M. To cover both natural and mathematical language, we train on the 50:50 mixture of OpenWebText (Gokaslan et al., 2019) and Proof-Pile-2 (Azerbayev et al., 2023). Surprisingly, we observe rapid transfer: within three days on $16$ H100 GPUs, the model generates variable-length sentences. We then instruction-fine-tune (IFT) this base FlexMDM to evaluate it on downstream tasks. See Appendix F for more details  [^3] For a fair comparison, since FlexMDM is not IFT-ed, we instruction fine-tune LLaDA-Base for the same number of epochs; this differs from Zhao et al. (2025).
Results. For comparison, we train FlexMDM and LLaDA-Base from the same number of IFT pairs. For math and code, respectively, we IFT on the GSM8K train (Cobbe et al. (2021); $\approx$8000 pairs) and the educational split of opc-sft-stage-2 (Huang et al., 2024) ($\approx$0.1M pairs), for which IFT-ed models are evaluated in GSM8K-test and HumanEval-infill (Bavarian et al., 2022) in zero-shot. Sampling is done by confidence-based sampling with sliding window. Notably, as the number of sampling steps increases, FlexMDM continues to improve, highlighting its strength in reasoning tasks given sufficient compute—whereas LLaDA’s performance remains flat. Although in this experiment we use IFT on task-specific pairs, we expect that, given sufficient compute, training on a much more diverse instruction–answer pairs will yield a more generalized model.
@@ Appendix A Related Works
Stochastic interpolant Albergo and Vanden-Eijnden [2022], Albergo et al. [2023a] is a general framework for building measure transport based generative models on continuous state space. While building off different philosophical grounds, it can be seen as equivalent to flow matcing Lipman et al. [2022], Liu et al. [2022]. Extensions of the interpolant has been proposed for conditional generation through data-dependent coupling Albergo et al. [2024] which we adopt for infilling, and for multi-task learning Negrel et al. [2025] by operator-based formulations.
@@
The most notable concurrent work is EditFlow Havasi et al. [2025] where the primary mathematical machinery that enabled their construction is a novel ”Flow Matching with Auxiliary Process”. We note that this can be seen to be mathematically equivalent to the notion of joint interpolant in this work. The main difference is philosophical: While EditFlow relies on the notion of probability path as its central mathematical construct, we write down the definition of two coupled random variables which only define the probability path implicitly. Such difference leads to a considerable difference in the interpolation used in the modeling, in the case of FlexMDM, our choice of the interpolation path enables any-order sampling.
@@ Appendix B Notation
In this section, we reiterate on the notations used in the main body and introduce auxiliary notations that are used in the proofs of the appenix.
@@
Let $\varepsilon$ denote the empty string, $\Sigma$ a vocabulary of words, $\mathbf{m}$ a special mask token. We write $\oplus$ for string concatenation, e.g., $ab\oplus bc=abbc$, $x^{i}$ the $i$-th element of $x$, $x[S]$ the string indexed by an index set, e.g. $abc[\{0,2\}]=ac$. To insert a token $v$ before position $i$ in string $x$, we write $x\,\triangleleft_{i}v$, e.g., to prepend a token $abc\,\triangleleft_{0}d=dabc$ and to append a token $abc\,\triangleleft_{4}d=abcd$. To replace the $i$-th token in $x$ with $v$, we write $x[x^{i}\leftarrow v]$. As much of the work involves masking, we write $x\subseteq y$ if $x$ can be constructed by partially masking $y$. We write $\mathrm{mask}(x),\mathrm{unmask}(x)\subseteq[\mathrm{len}(x)]$ for the set of indices corresponding to mask and clean tokens.
@@ Appendix C Discrete Stochastic Interpolants: Definitions and Propositions for Section 2
In the section C.1, we introduce a discrete analogue of the stochastic interpolant. To illustrate this framework, we reformulate the widely-used masked diffusion model for sequences of length $n$ within our setup. Briefly, masked diffusion defines the interpolation $p_{t}$ by progressively unmasking tokens in sentences drawn from the data distribution. At time $t=0$, all tokens are masked, so $p_{0}$ is a point mass at the fully masked sequence the $\mathbf{m}$ token repeated $n$ times. The transition rates driving the generative transport can be characterized as functions of per-token posterior probabilities conditioned on time $t$. These are typically learned by minimizing a variational objective in the form of a weighted cross-entropy loss.
@@ C.1 Discrete Stochastic Interpolant
To obtain the target rate matrix, discrete stochastic interpolant relies on an interpolating rate matrix that drives sample from a sample drawn from a sample from $p_{0}$ to a sample from $p_{1}$, defined as follows:
@@
With an interpolating rate of an interpolant, the target rate matrix can then be obtained trough Proposition C.1
@@
Writing $p_{t}$ a pmf of $x_{t}$ and $p_{t}(\cdot|x_{0},x_{1})$ a pmf of $x_{t}$ conditioned on $x_{0},x_{1}$. We further write $q(x_{0},x_{1})$ the joint pmf of $x_{0}$ and $x_{1}$ and $q_{t}(x_{0},x_{1}|x_{t})$ to tbe the joint pmf conditioned on $x_{t}$
It suffices to show $R_{t}$ satifies the Kolmogorov Forward Equation with pmf $p_{t}$ as follows:
@@
While written considerably differently, the framework is mathematically equivalent to discrete flow matching Gat et al. [2024]. The difference is only philosophical: discrete flow matching relies on the notion of a conditional probability path that the interpolating rate should induce, whereas we define such probability path only implicitly through the definition of the interpolant.
@@ C.2 The Masked Diffusion Interpolant
As a concrete example of the discrete stochastic interpolant, we reformulate the masked diffusion model and its learning under the framework. As masked diffusion starts from a point mass, we drop the dependence of $x_{0}$ in the writing.
@@
Let $x_{1}\sim p_{1}$ be a sentence of length $n$ drawn from the data and $\alpha_{t}$ a smooth unmasking schedule that interpolates from $\alpha_{t=0}=0$ to $\alpha_{t=1}=1$. Define the unmasking times $\{T^{i}\}_{i\in[1,\dots,n]}$ as:
$\forall i\in\{1,\dots,n\}:\quad T^{i}\sim\dot{\alpha_{t}}\;dt,$ (9)
@@
In words, at each time $t$, $x_{t}$ reveals a subset of the tokens from $x_{1}$, with each token $x_{1}^{i}$ independently unmasked at its associated time $T^{i}$.
@@
Let $p_{t}(\cdot|x_{1})$ as the pmf of $x_{t}$ conditioned on $x_{1}$, from the definition of the interpolant, we notice that:
$p_{t}(x|x_{1})=\prod_{i=0}^{\text{len}(x_{1})-1}(1-\alpha_{t})\mathbf{1}\{x^{i}=\mathbf{m}\}+\alpha_{t}\mathbf{1}\{x^{i}=x_{1}^{i}\}$
We verify that $K_{t}$ satifies the Kolmogorov Forward Equation (3) under the conditioned pmf as follows,
$\text{L.H.S}=\partial_{t}p_{t}(x|x_{1})$ $=\partial_{t}\prod_{i=0}^{\text{len}(x_{1})-1}(1-\alpha_{t})\mathbf{1}\{x^{i}=\mathbf{m}\}+\alpha_{t}\mathbf{1}\{x^{i}=x_{1}^{i}\}$
$=\sum_{i=0}^{\text{len}(x_{1})-1}\left(-\dot{\alpha_{t}}\mathbf{1}\{x^{i}=\mathbf{m}\}+\dot{\alpha_{t}}\mathbf{1}\{x^{i}=x_{1}^{i}\}\right)\cdot\prod_{j\neq i}(1-\alpha_{t})\mathbf{1}\{x^{j}=\mathbf{m}\}+\alpha_{t}\mathbf{1}\{x^{j}=x_{j}^{i}\}$
R.H.S $=\sum_{i=0}^{\text{len}(x_{t})-1}\mathbf{1}\{x^{i}=x_{1}^{i}\}\frac{\dot{\alpha_{t}}}{1-\alpha_{t}}p_{t}(x[x^{i}\leftarrow\mathbf{m}]|x_{1})-\mathbf{1}\{x^{i}=\mathbf{m}\}\frac{\dot{\alpha_{t}}}{1-\alpha_{t}}p_{t}(x|x_{1})$
@@
To learn an approximation to $R_{t}$, we now parameterize an approximate target rate of the form $\hat{R}_{t}(x,x[x^{i}\leftarrow v])=\frac{\dot{\alpha}_{t}}{1-\alpha_{t}}\,f_{\theta}(x,t)[i,v]$ where $f_{\theta}(x,t)[i,v]$ is a learned approximation to the posterior $\mathbb{P}(x_{1}^{i}=v\mid x_{t}=x)$.
@@
The loss functional defined as:
$L[\hat{R}_{t}]$ $=\int_{0}^{1}\mathbb{E}_{x_{1},x_{t}}\left[-\frac{\dot{\alpha}_{t}}{1-\alpha_{t}}\sum_{i=1}^{n}\mathbf{1}\{x_{t}^{i}=\mathbf{m}\}\log f_{\theta}(x_{t},t)[i,x_{1}^{i}]\right]\;dt,$ (13)
@@
$=\int_{0}^{1}\mathbb{E}_{x_{1},x_{t}}\left[-\frac{\dot{\alpha}_{t}}{1-\alpha_{t}}\sum_{i=1}^{n}\mathbf{1}\{x_{t}^{i}=\mathbf{m}\}\log f_{\theta}(x_{t},t)[i,x_{1}^{i}]\right]\;dt+\text{Const.}$
where the first line takes expectation of the Radon-Nikodym derivative between the two path measures. The statement of Radon-Nikodym derivative between two CTMCs can be found in the Appendix in Campbell et al. [2024]. A discrete-time equivalent derivation can also be found in Shaul et al. [2025].
@@
This technique is standard as shown in Vargas et al. [2025] for the case of path reversal based construction of diffusion generative models, and in Holderrieth et al. [2025b] for discrete diffusion. ∎

### v2

@@ 3 Variable Length Masked Diffusions: Training
$s_{t}\colon=\{i\in\{0,\dots,\mathrm{len}(x_{1})-1\}\mid T_{1}^{i}\leq t\},$
@@
$f_{\theta}$ resembles the familiar unmasking posterior from MDMs, whereas $g_{\theta}$ is new: it predicts how many tokens need to be inserted. Intuitively, modeling a variable-length $p_{1}$ is harder than the fixed-length setup of MDM–introducing an insertion expectation allows us to parameterize more complicated CTMC for FlexMDM; its rate matrix will appear soon in Proposition 2. To define the training loss, we set the boundary values of $s_{t}$ as $s_{t}[-1]:=-1$ and $s_{t}[\mathrm{len}(s_{t})]:=\mathrm{len}(x_{1})$, and let $\phi(x,y)=y-x\log y$ denote a scalar Bregman divergence.
@@ 5.1.1 Pretraining on text data
We first construct a training dataset from the raw OpenWebText corpus (Gokaslan et al., 2019), splitting each article into paragraphs to preserve semantic coherence and yield variable-length sequences. Models pretrained on this data, therefore, generate variable-length text.
Results. We train 175175M FlexMDM and MDM with a maximum sequence length $1024$ for 500500K iterations and batch size $1024$. Using the pretrained models, we vary the number of sampling steps and measure (a) generative perplexity as a proxy for text fluency, and (b) the induced length distribution. Figure 5(c) shows comparative generative perplexity for the two models, improving as sampling steps increase, indicating no fluency degradation for FlexMDM despite its more involved loss objective. Crucially, we observe that FlexMDM matches the true length distribution far more closely (Figure 5(b)): with only $256$ steps it tracks the ground truth distribution (red line, whereas MDM remains miscalibrated even at $1024$ steps (blue line).
@@ 5.1.2 Planning task
We further evaluate FlexMDM’s ability in a planning task in a discrete space. Motivated by an earlier study Janner et al. (2022) that investigated the ability of continuous diffusion in maze tasks, we design a grid-maze benchmark: the maze is fixed but unknown to the model, with a subset of cells invalid. Given a sequence of subgoal grids $(g_{1},\dots,g_{K})$, the model must connect this sequence without entering invalid cells (see Figure 5(a)). This subgoal structure aligns naturally with FlexMDM: starting from $(g_{1},\dots,g_{K})$, inference inserts mask tokens between subgoals and then unmasks to generate a feasible path. Theoretically, this can be seen as augmenting the base distribution to contain a data-dependent distribution Albergo et al. (2024). This is in stark contrast to MDM, where it must preassign each subgoal to a specific position, which is difficult to know a priori. We provide additional details in Appendix F.2.
@@ 5.2 Scaling up FlexMDM
In this section, we address FlexMDM’s scalability by scaling it to 8B parameters and observing notable improvements over an MDM baseline. We start from the observation that MDM and FlexMDM both share the unmasking posterior as a core component, suggesting effective task transfer from a pretrained MDM might be possible. To demonstrate this, concretely, we initialize from LLaDA-Base (Nie et al., 2025) and make the following modifications: (a) add time-embedding layers and a scalar head to model the insertion expectation; (b) attach LoRA adapters. Altogether, the resulting number of trainable parameters is $\approx$400M. To cover both natural and mathematical language, we train on the 50:50 mixture of OpenWebText (Gokaslan et al., 2019) and Proof-Pile-2 (Azerbayev et al., 2023). Surprisingly, we observe rapid transfer: within three days on $16$ H100 GPUs, the model generates variable-length sentences. We then instruction-fine-tune (IFT) this base FlexMDM to evaluate it on downstream tasks. See Appendix F for more details  [^3] For a fair comparison, since FlexMDM is not IFT-ed, we IFT LLaDA-Base, rather LLaDA-instruct, this differs from Zhao et al. (2025). We employ zero-shot evaluation, which also differs from Nie et al. (2024).
Results. For comparison, we train FlexMDM and LLaDA-Base from the same number of IFT pairs. For math and code, respectively, we IFT on the GSM8K train split(Cobbe et al. (2021); $\approx$8000 pairs) and the educational split of opc-sft-stage-2 (Huang et al., 2024) ($\approx$0.1M pairs), for which IFT-ed models are evaluated in the GSM8K test split and HumanEval-infill (single line) (Bavarian et al., 2022) in zero-shot. Sampling is done by confidence-based sampling with a sliding window. Notably, as the number of sampling steps increases, FlexMDM continues to improve, highlighting its strength in reasoning tasks given sufficient compute—whereas the IFT-ed LLaDA’s performance remains flat. Although in this experiment we use IFT on task-specific pairs, we expect that training on a much more diverse instruction–answer pairs with sufficient compute will yield a more generalized model.
@@ Appendix A Related Works
Stochastic interpolant  [Albergo and Vanden-Eijnden, 2022, Albergo et al., 2023a] is a general framework for building measure-transport based generative models on continuous state space. While building off different philosophical grounds, it can be seen as equivalent to flow matching [Lipman et al., 2022, Liu et al., 2022]. Extensions of the interpolant have been proposed for conditional generation through data-dependent coupling  [Albergo et al., 2024], which we adopt for infilling task design in Section 5.
@@
The most notable concurrent work is EditFlow [Havasi et al., 2025], where the primary mathematical machinery that enabled their construction is referred to as “Flow Matching with Auxiliary Process”. We note that this can be interpreted as mathematically equivalent to the notion of joint interpolant in this work, e.g., our Proposition D.1 is equivalent to Theorem 3.1 in Havasi et al. [2025].
The main differences are (1) the choice of interpolant and (2) the guarantee of any-order inference. Whereas EditFlow is built around an explicit probability path, we instead define a pair of coupled random variables that implicitly induce this path, leading to a different choice of intermediate. As discussed in Section 3, our choice of interpolant yields a distinct training objective for an unmasking posterior and an insertion-expectation term. Consequently, it enables the any-order inference guarantee established in Section 4.
@@ Appendix B Notation
In this section, we reiterate the notations used in the main body and introduce auxiliary notations that are used in the proofs of the appendix.
@@
Let $\varepsilon$ denote the empty string, $\Sigma$ a vocabulary of words, $\mathbf{m}$ a special mask token. We write $x^{i}$ the $i$-th element of $x$ with 0-baesd indexing, $x|_{S}$ the string indexed by an index set, e.g. $abc[\{0,2\}]=ac$. To insert a token $v$ before position $i$ in string $x$, we write $x\,\triangleleft_{i}v$, e.g., to prepend a token $abc\,\triangleleft_{0}d=dabc$ and to append a token $abc\,\triangleleft_{3}d=abcd$. To replace the $i$-th token in $x$ with $v$, we write $x[x^{i}\leftarrow v]$. As much of the work involves masking, we write $x\subseteq y$ if $x$ can be constructed by partially masking $y$. We write $\mathrm{mask}(x),\mathrm{unmask}(x)\subseteq[\mathrm{len}(x)]$ for the set of indices corresponding to mask and clean tokens.
@@ Appendix C Discrete Stochastic Interpolants: Definitions and Propositions for Section 2
In the section C.1, we introduce a discrete analogue of the stochastic interpolant. To illustrate this framework, we reformulate the widely used masked diffusion model for sequences of length $n$ within our setup. Briefly, masked diffusion defines the interpolation $p_{t}$ by progressively unmasking tokens in sentences drawn from the data distribution. At time $t=0$, all tokens are masked, so $p_{0}$ is a point mass at the fully masked sequence, the $\mathbf{m}$ token repeated $n$ times. The transition rates driving the generative transport can be characterized as functions of per-token posterior probabilities conditioned on time $t$. These are typically learned by minimizing a variational objective in the form of a weighted cross-entropy loss.
@@ C.1 Discrete Stochastic Interpolant
To obtain the target rate matrix, the discrete stochastic interpolant relies on an interpolating rate matrix that drives a sample from a sample drawn from a sample from $p_{0}$ to a sample from $p_{1}$, defined as follows:
@@
With an interpolating rate of an interpolant, the target rate matrix can then be obtained through Proposition C.1
@@
Writing $p_{t}$ a probability mass function (pmf) of $x_{t}$ and $p_{t}(\cdot|x_{0},x_{1})$ a pmf of $x_{t}$ conditioned on $x_{0},x_{1}$. We further write $q(x_{0},x_{1})$ the joint pmf of $x_{0}$ and $x_{1}$ and $q_{t}(x_{0},x_{1}|x_{t})$ to be the joint pmf conditioned on $x_{t}$.
It suffices to show $R_{t}$ satisfies the Kolmogorov Forward Equation with pmf $p_{t}$ as follows:
@@
While written considerably differently, the framework is mathematically equivalent to discrete flow matching Gat et al. [2024]. The difference is only philosophical: discrete flow matching relies on the notion of a conditional probability path that the interpolating rate should induce, whereas we define such a probability path only implicitly through the definition of the interpolant.
@@ C.2 The Masked Diffusion Interpolant
As a concrete example of the discrete stochastic interpolant, we reformulate the masked diffusion model and its learning in the framework. As masked diffusion starts from a point mass, we drop the dependence of $x_{0}$ in writing.
@@
Let $x_{1}\sim p_{1}$ be a sentence of length $n$ drawn from the data and $\alpha_{t}$ a smooth unmasking schedule that interpolates from $\alpha_{t=0}=0$ to $\alpha_{t=1}=1$. Define the unmasking times $\{T^{i}\}_{i\in[0,\dots,n-1]}$ as:
$\forall i\in\{0,\dots,n-1\}:\quad T^{i}\sim\dot{\alpha_{t}}\;dt,$ (9)
@@
In other words, at each time $t$, $x_{t}$ reveals a subset of the tokens of $x_{1}$, with each token $x_{1}^{i}$ independently unmasked at its associated time $T^{i}$.
@@
Let $p_{t}(\cdot|x_{1})$ as the pmf of $x_{t}$ conditioned on $x_{1}$. From the definition of the interpolant, we notice that:
$p_{t}(x|x_{1})=\prod_{i=0}^{\text{len}(x_{1})-1}\left[(1-\alpha_{t})\mathbf{1}\{x^{i}=\mathbf{m}\}+\alpha_{t}\mathbf{1}\{x^{i}=x_{1}^{i}\}\right]$
We verify that $K_{t}$ satisfies the Kolmogorov Forward Equation (3) under the conditioned pmf as follows,
L.H.S
$=\partial_{t}p_{t}(x|x_{1})$
$=\partial_{t}\left[\prod_{i=0}^{\text{len}(x_{1})-1}(1-\alpha_{t})\mathbf{1}\{x^{i}=\mathbf{m}\}+\alpha_{t}\mathbf{1}\{x^{i}=x_{1}^{i}\}\right]$
$=\sum_{i=0}^{\text{len}(x_{1})-1}\left(-\dot{\alpha_{t}}\mathbf{1}\{x^{i}=\mathbf{m}\}+\dot{\alpha_{t}}\mathbf{1}\{x^{i}=x_{1}^{i}\}\right)\cdot\prod_{j\neq i}(1-\alpha_{t})\mathbf{1}\{x^{j}=\mathbf{m}\}+\alpha_{t}\mathbf{1}\{x^{j}=x_{j}^{i}\},$
R.H.S
$=\sum_{i=0}^{\text{len}(x_{t})-1}\mathbf{1}\{x^{i}=x_{1}^{i}\}\frac{\dot{\alpha_{t}}}{1-\alpha_{t}}p_{t}(x[x^{i}\leftarrow\mathbf{m}]|x_{1})-\mathbf{1}\{x^{i}=\mathbf{m}\}\frac{\dot{\alpha_{t}}}{1-\alpha_{t}}p_{t}(x|x_{1})$
@@
To learn an approximation to $R_{t}$, we now parameterize an approximate target rate of the form $\hat{R}_{t}(x,x[x^{i}\leftarrow v]):=\frac{\dot{\alpha}_{t}}{1-\alpha_{t}}\,f_{\theta}(x,t)[i,v]$ where $f_{\theta}(x,t)[i,v]$ is a learned approximation to the posterior $\mathbb{P}(x_{1}^{i}=v\mid x_{t}=x)$.
@@
The loss function is defined as:
$L[\hat{R}_{t}]$ $=\int_{0}^{1}\mathbb{E}_{x_{1},x_{t}}\left[-\frac{\dot{\alpha}_{t}}{1-\alpha_{t}}\sum_{i=0}^{n-1}\mathbf{1}\{x_{t}^{i}=\mathbf{m}\}\log f_{\theta}(x_{t},t)[i,x_{1}^{i}]\right]\;dt,$ (13)
@@
$=\int_{0}^{1}\mathbb{E}_{x_{1},x_{t}}\left[-\frac{\dot{\alpha}_{t}}{1-\alpha_{t}}\sum_{i=0}^{n-1}\mathbf{1}\{x_{t}^{i}=\mathbf{m}\}\log f_{\theta}(x_{t},t)[i,x_{1}^{i}]\right]\;dt+\text{Const.}$
where the first line takes the expectation of the Radon-Nikodym derivative between the two path measures. The statement of Radon-Nikodym derivative between two CTMCs can be found in the Appendix in Campbell et al. [2024]. A discrete-time equivalent derivation can also be found in Shaul et al. [2025].
@@
This technique is standard, as shown in Vargas et al. [2025] for the case of path reversal-based construction of diffusion generative models, and in Holderrieth et al. [2025b] for discrete diffusion. ∎